A mass of 0.3634 g of an unknown acid, HA, is titrated with NaOH. If the acid reacts with 26.63 mL of 0.1195 M NaOH(aq), what is the molar mass of the acid?

A. 0.3634 g/mol
B. 3.041 g/mol
C. 0.1142 g/mol
D. 1.156×10-3 g/mol
E. 114.2 g/mol
1
Expert's answer
2019-03-11T08:17:40-0400

HA + NaOH = NaA + H2O

n(NaOH) = c * V = 0.1195 * 26.63 = 3,18 (mmol)

n(HA) = n(NaOH)

M(HA) = m / n = 0.3634 / 0.00318 = 114.2 (g/mol)


Answer - E, 114.2 g/mol